# REBEL: Reinforcement Learning via Regressing Relative Rewards

### v1

@@ 3 Understanding REBEL as an Adaptive Policy Gradient
We begin by recapping the foundational algorithms for policy optimization before situating REBEL within this space of techniques.
@@ 3.1 Adaptive Gradient Algorithms for Policy Optimization
In this section, we give a brief overview of three adaptive gradient algorithms: Mirror Descent (MD), Natural Policy Gradient (NPG), and Proximal Policy Optimization (PPO). We discuss why they are preferable to their non-adaptive counterparts (Gradient Descent (GD) and Policy Gradient (PG)) and the connections between them.
@@
where $F_{t}^{\dagger}$ is pseudo-inverse of $F_{t}$, and $\mathbb{E}_{x,y\sim\pi_{\theta_{t}}(\cdot|x)}\nabla\ln\pi_{\theta_{t}}(y|x)r(x,y)$ is the standard policy gradient (i.e. REINFORCE (Williams, 1992)). As mentioned above, this update procedure can be understood as performing gradient updates in the local geometry induced by the Fisher information matrix, which ensures that we are taking small steps in policy space rather than in parameter space. Conversely, unlike regular gradient descent methods (i.e., PG), NPG allows us to make large changes in the parameter space $\Theta$, as long as the resulting two policies are close to each other in terms of KL divergence. This property allows NPG to make more aggressive and adaptive updates in the parameter space of the policy as well as be invariant to linear transformations of the parameters. Theoretically, Agarwal et al. (2021a) show that NPG with softmax parameterization converges at the $1/T$ rate in a dimension-free manner, provably faster than the standard PG under the same setup. Empirically, the superior convergence speed of NPG compared to that of PG was observed in its original exploration (Kakade, 2001; Bagnell and Schneider, 2003), as well as in follow-up work like TRPO (Schulman et al., 2015a). Critically, while elegant in theory, NPG, unfortunately, does not scale to modern generative models due to the need for computing the Fisher matrix inverse either explicitly or implicitly via the Hessian-vector matrix product trick.
@@
A REBEL With a Cause. Our algorithm REBEL addresses the limitations of NPG (scalability) and PPO (lack of conservativity or adaptivity) from above. First, unlike NPG, it does not rely on the Fisher information matrix at all and can easily scale to modern LLM applications, yet (as we will discuss below) can be interpreted as a generalization of NPG. Second, in contrast to PPO, it doesn’t have unjustified heuristics and thus enjoys strong convergence and regret guarantees just like NPG.
@@ 3.2 Connections between REBEL and MD / NPG
We now sketch a series of connections between REBEL and the methods outlined above.
@@
Conditioned on Eq. 12 being true, a few lines of algebraic manipulation reveals that there must exist a function $c(x)$ which is independent of $y$, such that:
$\forall x,y:\quad\frac{1}{\eta}\ln\frac{\pi_{\theta_{t+1}}(y|x)}{\pi_{\theta_{t}}(y|x)}=r(x,y)+c(x).$
Taking an $\exp$ on both sides and re-arrange terms, we get:
$\forall x,y:\quad\pi_{\theta_{t+1}}(y|x)\propto\pi_{\theta_{t}}(y|x)\exp\left(\eta r(x,y)\right).$
In other words, under the strong assumption that least square regression returns a point-wise accurate estimator (i.e., Eq. 12), we see the REBEL recovers the exact MD update, which gives it (a) a fast $1/T$ convergence rate (Shani et al., 2020; Agarwal et al., 2021a), (b) conservativity, i.e., $\max_{x}\text{KL}(\pi_{t+1}(\cdot|x)||\pi_{t}(\cdot|x))$ is bounded as long as $\max_{x,y}|r(x,y)|$ is bounded, and (c) monotonic policy improvement via the NPG standard analysis (Agarwal et al., 2021a).
NPG is Approximate REBEL with Gauss-Newton Updates. We provide another interpretation of REBEL by showing that NPG (Eq. 10) can be understood as a special case of REBEL where the least square problem in Eq. 9 is approximately solved via a single iteration of the Gauss-Newton algorithm.
As for any application of Gauss-Newton, we start by approximating our predictor $\frac{1}{\eta}\ln\pi_{\theta}(y|x)/\pi_{\theta_{t}}(y|x)$ by its first order Taylor expansion at $\theta_{t}$:
$\frac{1}{\eta}\left(\ln\pi_{\theta}(y|x)-\ln\pi_{\theta_{t}}(y|x)\right)\approx\frac{1}{\eta}\nabla_{\theta}\ln\pi_{\theta_{t}}(y|x)^{\top}(\theta-\theta_{t}),$
where $\approx$ indicates that we ignore higher order terms in the expansion. If we $\delta:=\theta-\theta_{t}$ and replace $\frac{1}{\eta}\left(\ln\pi_{\theta}(y|x)-\ln\pi_{\theta_{t}}(y|x)\right)$ by its above first order approximation in Eq. 9, we arrive at the following quadratic form:
$\min_{\delta}\mathbb{E}_{x\sim\rho,y\sim\pi_{\theta_{t}}(\cdot|x),y^{\prime}\sim\mu(\cdot|x)}\left(\frac{1}{\eta}\left(\nabla_{\theta}\ln\pi_{\theta_{t}}(y|x)-\nabla_{\theta}\ln\pi_{\theta_{t}}(y^{\prime}|x)\right)^{\top}\delta-\left(r(x,y)-r(x,y^{\prime})\right)\right)^{2}.$ (13)
Further simplifying notation, we denote the uniform mixture of $\pi_{t}$ and $\mu$ as $\pi_{mix}(\cdot|x):=(\pi_{t}(\cdot|x)+\mu(\cdot|x))/2$ and the Fisher information matrix $F_{t}$ averaged under said mixture as:
$F_{t}=\mathbb{E}_{x\sim\rho,y\sim\pi_{mix}(\cdot|x)}\left[\nabla_{\theta}\ln\pi_{\theta_{t}}(y|x)\left(\nabla_{\theta}\ln\pi_{\theta_{t}}(y|x)\right)^{\top}\right].$
Solving the above least square regression to obtain a minimum norm solution, we have the following claim.
@@
The minimum norm minimizer $\delta^{\star}$ of the least squares problem in Eq. 13 recovers an advantage-based variant of the NPG update:
$\delta^{\star}:=\eta F_{t}^{\dagger}\left(\mathbb{E}_{x\sim\rho,y\sim\pi_{mix}(\cdot|x)}\nabla_{\theta}\ln\pi_{\theta_{t}}(y|x)[A^{\pi_{t}}(x,y)]\right),$
where $F_{t}^{\dagger}$ is pseudo-inverse of $F_{t}$, and the advantage is defined as $A^{\pi_{t}}(x,y):=r(x,y)-\mathbb{E}_{y^{\prime}\sim\pi_{t}(\cdot|x)}r(x,y)$.
The proof of this claim is deferred to Appendix A. Observe that in REBEL, we never explicitly compute the advantage $A^{\pi_{t}}$. However, applying Gauss-Newton to our objective leads to an advantage-based NPG (rather than the traditional $Q$-function based NPG, e.g., Q-NPG from Agarwal et al. (2021a, 2019)) which indicates that predicting reward difference has an implicit variance reduction effect, as by definition, an advantage function includes a value function baseline.  [^1] Note that the original form of NPG is on-policy (Kakade, 2001; Sutton et al., 1999), i.e., the expectations under $\pi_{t}$. Our formulation is more general: when set $\mu=\pi_{t}$, a Gauss-Newton step will recover the original on-policy form of NPG from Kakade (2001); Sutton et al. (1999). More recent works have extended NPG beyond on-policy (e.g., Agarwal et al. (2021a, 2020)).
3.3 Extending REBEL to General Preferences
3.3.1 A Game-Theoretic Perspective on Learning from Preferences
This implies that $(\pi_{\mathsf{MW}},\pi_{\mathsf{MW}})$ is a Nash Equilibrium (Wang et al., 2023b; Munos et al., 2023; Swamy et al., 2024; Ye et al., 2024). As is standard in game solving, our objective is to obtain an $\epsilon$-approximate MW $\widehat{\pi}$ measured by the duality gap (DG):
3.3.2 Self-Play Preference Optimization (SPO) with REBEL as Base Learner
@@ 4 Theoretical Analysis
In the previous section, we interpret REBEL as the exact MD and show its convergence by assuming that least square regression always returns a predictor that is accurate everywhere. While such an explanation is simple and has also been used in prior work, point-wise out-of-distribution generalization is an extremely strong condition and is significantly beyond what a standard supervised learning method can promise. In this section, we significantly relax this condition via a reduction-based analysis:
As long as we can solve the regression problems well in an in-distribution manner, REBEL can compete against any policy covered by the training data distributions.
Formally, we assume the following generalization condition holds on the regressors we find.
@@
Over $T$ iterations, assume that for all $t$, we have:
@@
for some $\epsilon$.
Intuitively, this assumption is saying that there is a function in our class of regressors that is able to accurately fit the difference of rewards. Recall that our class of regressors is isomorphic to our policy class. Therefore, as long as our class of policies is expressive, we would expect this assumption to hold with small $\epsilon$. For all domains we consider, our policy class is a flexible set of generative models (e.g. Transformer-based LLMs or diffusion models). Thus, we believe it is reasonable to believe this assumption holds in practice – see Figure 6 in Appendix G for empirical evidence of this point and Example 1 for more discussion.
@@
We say $\mu$ covers $\pi$ if $C_{\mu\to\pi}<+\infty$.
Our goal is to bound the regret between our learned policies and an arbitrary comparator $\pi^{*}$ (e.g. the optimal policy if it is covered by $\mu$) using $\epsilon$ and the concentrability coefficient defined in Eq. 15. The following theorem formally states the regret bound of our algorithm.
@@
Here the $O$-notation hides problem-dependent constants that are independent of $\epsilon,C_{\mu\to\pi^{*}},T$.
The above theorem shows a reduction from RL to supervised learning — as long as supervised learning works (i.e., $\epsilon$ is small), then REBEL can compete against any policy $\pi^{*}$ that is covered by the base data distribution $\mu$. In the regret bound, the $1/\sqrt{T}$ comes from Mirror Descent style update, and $C_{\mu\to\pi^{*}}\epsilon$ captures the cost of distribution shift: we train our regressors under distribution $\pi_{t}$ and $\mu$, but we want the learned regressor to predict well under $\pi^{*}$. Similar to the NPG analysis from Agarwal et al. (2021a), we now have a slower convergence rate $1/\sqrt{T}$, which is due to the fact that we have approximation error from learning. Such an agnostic regret bound — being able compete against any policy that is covered by training distributions – is the strongest type of agnostic learning results known in the RL literature, matching the best of what has appeared in prior policy optimization work including PSDP (Bagnell et al., 2003), CPI (Kakade and Langford, 2002), NPG (Agarwal et al., 2021a), and PC-PG (Agarwal et al., 2020). While in this work, we use the simplest and most intuitive definition of coverage – the density ratio-based definition in Eq. 15 – extension to more general ones such as transfer error (Agarwal et al., 2020, 2021a) or concentrability coefficients that incorporate function class (e.g., Song et al. (2023b)) is straightforward. We defer the proof of the above theorem and the detailed constants that we omitted in the $O$ notation to Appendix B.
4.1 Extension to General Preferences
Extending the above analysis to the general preference case is straightforward except that it requires a stronger coverage condition. This is because we want to find a Nash Equilibrium, which requires a comparison between the learned policy against all the other policies. Results from the Markov Game literature (Cui and Du, 2022b; Zhong et al., 2022; Cui and Du, 2022a; Xiong et al., 2023) and Cui and Du (2022b) have shown that the standard single policy coverage condition used in single-player optimization is provably not sufficient. In particular, they propose using a notion of unilateral concentrability for efficient learning, which can be defined as
Over $T$ iterations, assume that for all $t$, we have:
for some $\epsilon$.
Theorem 2.
We defer the proof to Appendix C. Note that the coverage condition here is much stronger than the single policy coverage condition in the RL setting. We conjecture that this is the cost one has to pay by moving to the more general preference setting and leaving the investigation of the necessarily coverage condition for future work.
@@ 5 Experiments
The implementation of REBEL follows Algorithm 1. In each iteration, REBEL collects a dataset $\mathcal{D}_{t}=\{x,y,y^{\prime}\}$, where $x\sim\rho,y\sim\pi_{t}(\cdot|x),y^{\prime}\sim\mu(\cdot|x)$. Subsequently, REBEL optimizes the least squares regression problem in Eq. 9 through gradient descent with AdamW (Loshchilov and Hutter, 2017). We choose $\mu=\pi_{t}$ such that both $y$ and $y^{\prime}$ are generated by the current policy. We empirically assess REBEL’s performance on both natural language generation and text-guided image generation.
5.1 Natural Language Generation
Baselines:
We compare REBEL with baseline RL algorithms, PPO (Schulman et al., 2017), Direct Preference Optimization (DPO) (Rafailov et al., 2023), and REINFORCE (Williams, 1992) and its multi-sample extension, REINFORCE Leave-One-Out (RLOO) (Kool et al., 2019). The REINFORCE method is implemented with a moving average baseline of the reward. We include two variants of RLOO with two ($k=2$) and four ($k=4$) generations per prompt.
Dataset:
We use the TL;DR summarization dataset (Stiennon et al., 2020) [^2] Dataset available at https://github.com/openai/summarize-from-feedback to train the model to generate summaries of Reddit posts based on human preference data. The dataset comprises human reference summaries and preference data. Following prior work (Stiennon et al., 2020; Rafailov et al., 2023; Ahmadian et al., 2024), we train the DPO baseline on the preference dataset, while conducting online RL (PPO, RLOO, REBEL) on the human reference dataset. We set the maximum context length to 512 and the maximum generation length to 53 to ensure all references in the dataset can be generated. Additional dataset details are in Appendix D.1.
Models:
We include results with three different model sizes: 1.4B, 2.8B, and 6.9B. Each model is trained with a supervised fine-tuned (SFT) model and/or a reward model (RM) of the same size. For SFT models, we train a Pythia 1.4B (Biderman et al., 2023) [^3] HuggingFace Model Card: EleutherAI/pythia-1.4b-deduped model for 1 epoch over the dataset with human references as labels, and use the existing fine-tuned 2.8B [^4] HuggingFace Model Card: vwxyzjn/EleutherAI_pythia-2.8b-deduped__sft__tldr and 6.9B [^5] HuggingFace Model Card: vwxyzjn/EleutherAI_pythia-6.9b-deduped__sft__tldr models. For reward models, we train a Pythia 1.4B parameter model for 1 epoch over the preference dataset and use the existing reward models with 2.8B [^6] HuggingFace Model Card: vwxyzjn/EleutherAI_pythia-2.8b-deduped__reward__tldr and 6.9B [^7] HuggingFace Model Card: vwxyzjn/EleutherAI_pythia-6.9b-deduped__reward__tldr parameters. For both REBEL and baseline methods using 1.4B and 2.8B parameters, we trained the policy and/or the critic using low-rank adapters (LoRA) (Hu et al., 2022) on top of our SFT and/or reward model respectively. For the 6.9B models, we perform full-parameter training. More details about the hyperparameters are described in Appendix D.2.
Evaluation:

### v2

@@ 3 Understanding REBEL as an Adaptive Policy Gradient
In this section, we interpret REBEL as an adaptive policy gradient method to illuminate the relationship to past techniques. We start by introducing algorithms such as Mirror Descent, NPG, and PPO, followed by illustrating why REBEL addresses the limitations of these past algorithms.
@@ 3.1 Adaptive Gradient Algorithms for Policy Optimization
where $F_{t}^{\dagger}$ is pseudo-inverse of $F_{t}$, and $\mathbb{E}_{x,y\sim\pi_{\theta_{t}}(\cdot|x)}\nabla\ln\pi_{\theta_{t}}(y|x)r(x,y)$ is the standard policy gradient (i.e. REINFORCE (Williams, 1992)). As mentioned above, this update procedure can be understood as performing gradient updates in the local geometry induced by the Fisher information matrix, which ensures that we are taking small steps in policy space rather than in parameter space. Conversely, unlike regular gradient descent methods (i.e., PG), NPG allows us to make large changes in the parameter space $\Theta$, as long as the resulting two policies are approximately close to each other in terms of KL divergence. This property allows NPG to make more aggressive and adaptive updates in the parameter space of the policy as well as be invariant to linear transformations of the parameters. Theoretically, Agarwal et al. (2021) show that NPG with softmax parameterization converges at the $1/T$ rate in a dimension-free manner, provably faster than the standard PG under the same setup. Empirically, the superior convergence speed of NPG compared to that of PG was observed in its original exploration (Kakade, 2001; Bagnell and Schneider, 2003), as well as in follow-up work like TRPO (Schulman et al., 2015). Critically, while elegant in theory, NPG, unfortunately, does not scale to modern generative models due to the need for computing the Fisher matrix inverse either explicitly or implicitly via the Hessian-vector matrix product trick.
@@ 3.2 Connections between REBEL and MD / NPG
Conditioned on Eq. 8 being true, a few lines of algebraic manipulation reveal that there must exist a function $c(x)$ which is independent of $y$, such that $\forall x,y:\frac{1}{\eta}\ln\frac{\pi_{\theta_{t+1}}(y|x)}{\pi_{\theta_{t}}(y|x)}=r(x,y)+c(x)$. Taking an $\exp$ on both sides and re-arrange terms, we get $\forall x,y:\pi_{\theta_{t+1}}(y|x)\propto\pi_{\theta_{t}}(y|x)\exp\left(\eta r(x,y)\right)$. In other words, under the strong assumption that least square regression returns a point-wise accurate estimator (i.e., Eq. 8), we see the REBEL recovers the exact MD update, which gives it (a) a fast $1/T$ convergence rate (Shani et al., 2020; Agarwal et al., 2021), (b) conservativity, i.e., $\max_{x}\text{KL}(\pi_{t+1}(\cdot|x)||\pi_{t}(\cdot|x))$ is bounded as long as $\max_{x,y}|r(x,y)|$ is bounded, and (c) monotonic policy improvement via the NPG standard analysis (Agarwal et al., 2021).
NPG is Approximate REBEL with Gauss-Newton Updates. We provide another interpretation of REBEL by showing that NPG (Eq. 6) can be understood as a special case of REBEL where the least square problem in Eq. 1 is approximately solved via a single iteration of the Gauss-Newton algorithm. We start by approximating our predictor $\frac{1}{\eta}\ln\pi_{\theta}(y|x)/\pi_{\theta_{t}}(y|x)$ by its first order Taylor expansion at $\theta_{t}$: $\frac{1}{\eta}\left(\ln\pi_{\theta}(y|x)-\ln\pi_{\theta_{t}}(y|x)\right)\approx\frac{1}{\eta}\nabla_{\theta}\ln\pi_{\theta_{t}}(y|x)^{\top}(\theta-\theta_{t})$, where $\approx$ indicates that we ignore higher order terms in the expansion. Define $\delta:=\theta-\theta_{t}$ and replace $\frac{1}{\eta}\left(\ln\pi_{\theta}(y|x)-\ln\pi_{\theta_{t}}(y|x)\right)$ by its first order approximation in Eq. 1. Then, we have :
$\min_{\delta}\mathbb{E}_{x\sim\rho,y\sim\pi_{\theta_{t}}(\cdot|x),y^{\prime}\sim\mu(\cdot|x)}\left(\frac{1}{\eta}\left(\nabla_{\theta}\ln\pi_{\theta_{t}}(y|x)-\nabla_{\theta}\ln\pi_{\theta_{t}}(y^{\prime}|x)\right)^{\top}\delta-\left(r(x,y)-r(x,y^{\prime})\right)\right)^{2}$ (9)
Further simplifying notation, we denote the uniform mixture of $\pi_{t}$ and $\mu$ as $\pi_{mix}(\cdot|x):=(\pi_{t}(\cdot|x)+\mu(\cdot|x))/2$ and the Fisher information matrix $F_{t}$ averaged under said mixture as $F_{t}=\mathbb{E}_{x\sim\rho,y\sim\pi_{mix}(\cdot|x)}\left[\nabla_{\theta}\ln\pi_{\theta_{t}}(y|x)\left(\nabla_{\theta}\ln\pi_{\theta_{t}}(y|x)\right)^{\top}\right]$. Solving the above least squares problem to obtain a minimum norm solution, we have the following result.
@@
The minimum norm minimizer $\delta^{\star}$ of the least squares problem in Eq. 9 recovers an advantage-based NPG update: $\delta^{\star}:=\eta F_{t}^{\dagger}\left(\mathbb{E}_{x\sim\rho,y\sim\pi_{mix}(\cdot|x)}\nabla_{\theta}\ln\pi_{\theta_{t}}(y|x)[A^{\pi_{t}}(x,y)]\right)$ where $F_{t}^{\dagger}$ is pseudo-inverse of $F_{t}$, and the advantage is defined as $A^{\pi_{t}}(x,y):=r(x,y)-\mathbb{E}_{y^{\prime}\sim\pi_{t}(\cdot|x)}r(x,y)$.
The proof of this claim is deferred to Appendix A.
The implicit variance reduction effect of REBEL We show that regressing to relative rewards has a variance reduction effect by extending the previous derivation on REBEL with Gauss-Newton update to the setting of finite data $\mathcal{D}=\{x_{n},y_{n},y_{n}^{\prime}\}_{n=1}^{N}$. Denote the unbiased estimate of the Fisher information matrix as $\hat{F}_{t}=\frac{1}{N}\sum_{n=1}^{N}\left[\nabla_{\theta}\ln\pi_{\theta_{t}}(y_{n}|x_{n})\left(\nabla_{\theta}\ln\pi_{\theta_{t}}(y_{n}|x_{n})\right)^{\top}\right]$ and have the following claim.
Claim 2.
The minimum norm minimizer $\delta^{\star}$ in Eq. 9 under finite setting has the form $\delta^{\star}:=\eta\hat{F}_{t}^{\dagger}\frac{1}{2N}\sum_{n}\left(\nabla\ln\pi_{\theta_{t}}(y_{n}|x_{n})(r(x_{n},y_{n})-r(x_{n},y^{\prime}_{n}))+\nabla\ln\pi_{\theta_{t}}(y^{\prime}_{n}|x_{n})(r(x_{n},y^{\prime}_{n})-r(x_{n},y_{n}))\right)$ where $\hat{F}_{t}^{\dagger}$ is pseudo-inverse of $\hat{F}_{t}$.
The proof of this claim is deferred to Appendix B. Looking at the gradient formulation $\nabla\ln\pi_{\theta_{t}}(y_{n}|x_{n})\left(r(x_{n},y_{n})-r(x_{n},y_{n}^{\prime})\right)$ in $\delta^{\star}$, we see that $r(x_{n},y^{\prime}_{n})$ serves as a baseline for variance reduction. Interestingly, this gradient formulation is similar to RLOO (REINFORCE with leave-one-out) (Kool et al., 2019). However, different from RLOO, we pre-condition this variance reduced policy gradient formulation via the Fisher information matrix, leading to better performance.
A REBEL With a Cause. Our algorithm REBEL addresses the limitations of NPG (scalability) and PPO (lack of conservativity or adaptivity) from above. First, unlike NPG, it does not rely on the Fisher Information Matrix at all and can easily scale to modern LLM and image generation applications, yet can be interpreted as a generalization of NPG. Second, in contrast to PPO, it doesn’t have unjustified heuristics and thus enjoys strong convergence and regret guarantees just like NPG. Building on Swamy et al. (2024), we also show how to extend REBEL to preference-based settings without assuming transitivity in Appendix C.
@@ 4 Theoretical Analysis
In the previous section, we interpret REBEL as exact MD and show its convergence by assuming that least square regression always returns a predictor that is accurate everywhere. While such an explanation is simple and has also been used in prior work (Calandriello et al., 2024; Rosset et al., 2024), point-wise out-of-distribution generalization is an extremely strong condition and is significantly beyond what a standard supervised learning method can promise. In this section, we substantially relax this condition via a reduction-based analysis: As long as we can solve the regression problems well in an in-distribution manner, REBEL can compete against any policy covered by the training data distributions. Formally, we assume the following generalization condition holds on the regressors we find.
@@
Over $T$ iterations, assume that for all $t$, we have the following for some $\epsilon$:
@@
Intuitively, this assumption is saying that there is a function in our class of regressors that is able to accurately fit the difference of rewards. Recall that our class of regressors is isomorphic to our policy class. Therefore, as long as our class of policies is expressive, we would expect this assumption to hold with small $\epsilon$. For all domains we consider, our policy class is a flexible set of generative models (e.g. Transformer-based LLMs or diffusion models). Thus, we believe it is reasonable to believe this assumption holds in practice – see Figure 6 in Appendix J for empirical evidence of this point and Example 1 for more discussion.
@@
We say $\mu$ covers $\pi$ if $C_{\mu\to\pi}<+\infty$. Our goal is to bound the regret between our learned policies and an arbitrary comparator $\pi^{*}$ (e.g. the optimal policy if it is covered by $\mu$) using $\epsilon$ and the concentrability coefficient defined in Eq. 10. The following theorem formally states the regret bound of our algorithm.
@@
The above theorem shows a reduction from RL to supervised learning — as long as supervised learning works (i.e., $\epsilon$ is small), then REBEL can compete against any policy $\pi^{*}$ that is covered by the base data distribution $\mu$. In the regret bound, the $1/\sqrt{T}$ comes from Mirror Descent style update, and $C_{\mu\to\pi^{*}}\epsilon$ captures the cost of distribution shift: we train our regressors under distribution $\pi_{t}$ and $\mu$, but we want the learned regressor to predict well under $\pi^{*}$. Similar to the NPG analysis from Agarwal et al. (2021), we now have a slower convergence rate $1/\sqrt{T}$, due to the fact that we have approximation error from learning. Such an agnostic regret bound — being able to compete against any policy that is covered by training distributions — is the strongest type of agnostic learning results known in the RL literature, matching the best of what has appeared in prior policy optimization work including PSDP (Bagnell et al., 2003), CPI (Kakade and Langford, 2002), NPG (Agarwal et al., 2021), and PC-PG (Agarwal et al., 2020). While in this work we use the simplest and most intuitive definition of coverage – the density ratio-based definition in Eq. 10 – extension to more general ones such as transfer error (Agarwal et al., 2020, 2021) or concentrability coefficients that incorporate the function class (e.g., Song et al. (2023)) is straightforward. We defer the proof of the above theorem and the detailed constants that we omitted in the $O$ notation to Appendix D. We include an extension of the above analysis to the general preference setting in Appendix E.
@@ 5 Experiments
Our implementation of REBEL closely follows the psuedocode in Algorithm 1. In each iteration, REBEL collects a dataset $\mathcal{D}_{t}=\{x,y,y^{\prime}\}$, where $x\sim\rho,y\sim\pi_{t}(\cdot|x),y^{\prime}\sim\mu(\cdot|x)$. Subsequently, REBEL optimizes the least squares regression problem in Eq. 1 through gradient descent with AdamW (Loshchilov and Hutter, 2017). We choose $\mu=\pi_{t}$ such that both $y$ and $y^{\prime}$ are generated by the current policy. We empirically assess REBEL’s performance on both natural language generation and text-guided image generation. Additional experiment details are in Appendix F.
5.1 Summarization
Task. We use the TL;DR dataset (Stiennon et al., 2020) where $x$ is a forum post from Reddit and $y$ is a summary generated by the policy. The dataset comprises human reference summaries and preference data. We compare REBEL with baseline RL algorithms, REINFORCE (Williams, 1992) and its multi-sample extension, REINFORCE Leave-One-Out (RLOO) (Kool et al., 2019), PPO (Schulman et al., 2017), Direct Preference Optimization (DPO) (Rafailov et al., 2023), and Iterative DPO (Guo et al., 2024). Our implementation of Iterative DPO replaces our square regression objective with the DPO objective where the binary preference labels are obtained based on the reward difference. The implementation detail of the baseline methods is provided in Appendix F.1.3. Following prior work (Stiennon et al., 2020; Rafailov et al., 2023; Ahmadian et al., 2024), we train DPO on the preference dataset, while conducting online RL (RLOO, PPO, Iterative DPO, REBEL) on the human reference dataset. We include results with three different model sizes: 1.4B, 2.8B, and 6.9B based on the pre-trained models from Pythia (Biderman et al., 2023). Each model is trained from a supervised fine-tuned (SFT) model using a reward model (RM) of the same size.